OpenStudy (anonymous):

Mr. Barns is moving merchandise into a new store. The merchandise is packed in box with these dimensions:length 18in, width 8in, and a height of 12in. Mr. Barns has 84 cubic ft. of cargo area in his truck. How many boxes can he fit in the cargo area.

OpenStudy (jlastino):

You first convert the inches into ft. 18/12 = 1.5ft 8/12 =2/3 12/12=1 Then find the volume of the box = lwh (1.5)(2/3)(1) Then divide 84 by the volume of the box to find how many he can fit into the cargo area